The computer-science aptitude score, x, and the achievement score, y (measured by a comprehensive final), were measured for 20 students in a beginning computer-science course. The results were as follows.

x 3 15 19 14 21 22 14 21 20 15 17 18 7 5 4 19 17 12 18 13
y 19 19 24 36 27 26 25 28 17 27 21 24 18 18 14 28 21 22 20 21

(a) Find the equation of the line of best fit. (Round your slope to four decimal places and your y-intercept to three decimal places.)
= _______+_______  x

(b) Find se2. (Round your answer to two decimal places.)

The computer-science aptitude score, x, and the achievement score, y (measured by a comprehensive final), were measured for 10 students in a beginning computer-science course. The results were as follows. Calculate the estimated standard error of regression, sb1. (Give your answer correct to two decimal places.)

x 4 16 20 13 22 21 15 20 19 16
y 18 18 23 35 28 27 26 29 16 26

_______________
Answer #1
x y (x-xbar)^2 (y-ybar)^2 (x-xbar)*(y-ybar)
3 19 136.89 14.0625 43.875
15 19 0.09 14.0625 -1.125
19 24 18.49 1.5625 5.375
14 36 0.49 175.5625 -9.275
21 27 39.69 18.0625 26.775
22 26 53.29 10.5625 23.725
14 25 0.49 5.0625 -1.575
21 28 39.69 27.5625 33.075
20 17 28.09 33.0625 -30.475
15 27 0.09 18.0625 1.275
17 21 5.29 3.0625 -4.025
18 24 10.89 1.5625 4.125
7 18 59.29 22.5625 36.575
5 18 94.09 22.5625 46.075
4 14 114.49 76.5625 93.625
19 28 18.49 27.5625 22.575
17 21 5.29 3.0625 -4.025
12 22 7.29 0.5625 2.025
18 20 10.89 7.5625 -9.075
13 21 2.89 3.0625 2.975
sum 294 455 646.2 485.75 282.5
mean 14.7 22.75 sxx syy sxy
slope=b1=sxy/sxx 0.437171154
intercept=b0=ybar-(slope*xbar) 16.32358403

# Regression equation is

y^=bo+b1*x

y^=16.324+0.437*x

Ansb:

x y (x-xbar)^2 (y-ybar)^2 (x-xbar)*(y-ybar)
4 18 158.76 43.56 83.16
16 18 0.36 43.56 3.96
20 23 11.56 2.56 -5.44
13 35 12.96 108.16 -37.44
22 28 29.16 11.56 18.36
21 27 19.36 5.76 10.56
15 26 2.56 1.96 -2.24
20 29 11.56 19.36 14.96
19 16 5.76 73.96 -20.64
16 26 0.36 1.96 -0.84
sum 166 246 252.4 312.4 64.4
mean 16.6 24.6 sxx syy sxy
slope=b1=sxy/sxx 0.255150555
intercept=b0=ybar-(slope*xbar) 20.36450079
SST SYY 312.4
SSR sxy^2/sxx 16.431696
SSE syy-sxy^2/sxx 295.9683
r^2 SSR/SST 0.0525983
error variance s^2 SSE/(n-2) 98.656101
S^2b1 s^2/sxx 0.390872
standard error b1=se(b1) sqrt(s^2b1) 0.6251976

#the estimated standard error of regression, sb1=0.63